graph the line that goes through the point (3,2) and has slope 2/5.
Answer by mananth(16946)

m= 2/5

Plug value of the slope and point (3 ,2) in
Y =mx+ b
2.00 = 1.2 + b
b= 2 1.2
b= 0.8
So the equation will be
Y = 2/5 x + 0.8